OpenStudy (anonymous):

sin 195°, without using tables.. Help please

OpenStudy (anonymous):

just type it in your calculator!

OpenStudy (angela210793):

sin195=sin(180+15)=-sin15

OpenStudy (anonymous):

We cant use calculators

OpenStudy (anonymous):

its \[-\sqrt{6}+\sqrt{2}/4 \]

hero (hero):

sin(225 - 30)

hero (hero):

sin(225) - sin(30) sorry

hero (hero):

Okay, I forgot. It's one of those...

OpenStudy (angela210793):

Hero aren't u supposed to express it as son(180+x) or smth 90+smth

hero (hero):

It doesn't have to have "180" in it , no

OpenStudy (angela210793):

well u mist express it as smth familiar ...wht sin225?

OpenStudy (anonymous):

√2/4(1-√3) this is in our workbook as result

hero (hero):

I mean -sqrt2/2

hero (hero):

Your workbook sucks

OpenStudy (anonymous):

SO please how can i get this result √2/4(1-√3) withou using calculator or tables? Yes, its pellet :(

hero (hero):

It's supposed to be something like what mfleisch got

OpenStudy (angela210793):

I got ur point Hero,sorry :):)

OpenStudy (anonymous):

I sont know how can they get tthis result √2/4(1-√3) ..

hero (hero):

It's some retarded long way of doing something very simple

OpenStudy (anonymous):

So, if i do sin(225) - sin(30) = 1.2 its right result?

OpenStudy (anonymous):

That is not correct You need to use this formula: sin(A-B)=sin A cos B - cos A sin B

OpenStudy (anonymous):

sin225=-sqrt2 /2 sin 30=0.5

OpenStudy (anonymous):

cos 225=-sqrt2 /2 cos 30=sqrt3 /2

OpenStudy (anonymous):

I think you have everything to solve it now

OpenStudy (anonymous):

ok i will try t, thanks :)

hero (hero):

Sorry, haven't done these in a while and I didn't memorize the formulas

hero (hero):

On my calculator, I got the same answer as in your book so I figured I must have been doing something wrong

hero (hero):

However, the answer is still equivalent to mfleisch's result. The only difference is, your book factored out sqrt{2}

OpenStudy (anonymous):

ok, thanks guys ;)